A boat moving downstream (with current in the river in 4min and 7s to cover a measeured mile according to marks on shore and 7min 30s to cover the same distance moving upstream (against the current) If the boats velocity relative to the water is 11 km/h. What is the velocity of the current in the river.


1
Expert's answer
2021-05-04T12:08:08-0400

"11\ km/h=3.056\ m/s"

"4\ min\ 7\ s=247\ s"

"7\ min\ 30\ s=450\ s"


"(3.056+v)\cdot247=(3.056-v)\cdot450\to v=0.89\ (m/s)" . Answer